OpenStudy (anonymous):

How to solve sin8x+sin4x=0. I think you use sum to product formula but I'm not sure. Also, there should be sixteen answers i believe..

OpenStudy (mertsj):

Sin(8x)=sin(4x+4x)

OpenStudy (anonymous):

sin8x+sin4x=0 sin(2(4x))+sin(2(2x))=0

OpenStudy (anonymous):

Use double angle formula sin2x = 2sinxcosx

OpenStudy (anonymous):

sin(2(4x))+sin(2(2x))=0 2sin(4x)cos(4x)+2sin(2x)cos(2x)=0

OpenStudy (anonymous):

can you further reduce that?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

2sin(2 * 2x)cos(2 * 2x)+2sin(2x)cos(2x)=0

OpenStudy (anonymous):

2[sin(2 * 2x)cos(2 * 2x)+sin(2x)cos(2x)]=0

OpenStudy (anonymous):

how did you get to that step? what equation?

OpenStudy (anonymous):

To simplify, I used the double angle identity for sin. Sin(2x) = 2sinxcosx. In each situation, I set "x" as either "4x" or "2x".

OpenStudy (anonymous):

2[sin(2 * 2x)cos(2 * 2x)+sin(2x)cos(2x)]=0 Divide both sides by 2. sin(2 * 2x)cos(2 * 2x) + sin(2x)cos(2x) = 0 2sin(2x)cos(2x)cos(2 *2x) + sin(2x)cos(2x) = 0 Factor sin(2x)cos(2x) * [2cos(2*2x) + 1] = 0 Use cos double angle identity cos(2x) = cos^x-sin^x sin(2x)cos(2x) * [2cos^x - 2sin^x + 1] = 0

OpenStudy (anonymous):

It get's quite complicated, to be honest.

OpenStudy (anonymous):

from there what terms would i set equal to zero to solve for my answer in radians?

OpenStudy (anonymous):

Keep in mind, the method I'm using here is the COMPLICATED WAY. Simpler way: sin(8x)+sin(4x)=0 sin(8x)=-sin(4x) sin(8x)=sin(-4x) Solve for x.

OpenStudy (anonymous):

how would i solve for x from there?(the easy way)

OpenStudy (anonymous):

because we are trying to not use calculators